Find the slope and y-intercept of the line that passes through(-4,-8) &(-3,-2)
Please show work
I would really appreciate any help I can get with this problem, Ijust dont understand this slope stuff at all. Please help me!!!!

Found 2 solutions by Emsley, Alan3354:
Answer by Emsley(35)

The slope (really called gradient)= the difference of the y coordinates divided by the difference of the x coordinates
So in this case it is (-2)-(-8) divided by (-3)-(-4) This gives 6/1=1

Answer by Alan3354(30982)

Find the slope and y-intercept of the line that passes through(-4,-8) &(-3,-2)
Please show work
-----------
Find the slope first.
Slope, m = diffy/diffx
m = 6/1 = 6
-------------
Use y = mx + b and the point to find b, the y-intercept
-8 = 6*(-4) + b
b = 16 or (0,16)
-------